Oriole Company is considering buying a new farm that it plans to operate for 10 years. The farm will require an initial investment of $12.10 million. This investment will consist of $2.00 million for land and $10.10 million for trucks and other equipment. The land, all trucks, and all other equipment are expected to be sold at the end of 10 years for a price of $5.00 million, which is $2.40 million above book value. The farm is expected to produce revenue of $2.05 million each year, and annual cash flow from operations equals $1.85 million. The marginal tax rate is 35 percent, and the appropriate discount rate is 10 percent. Calculate the NPV of this investment. (Do not round factor values. Round final answer to 2 decimal places, e.g. 15.25.)

NPV = $
Answer #1

Formulas used:-

Depriciation=(12100000-2600000)/10
Profit before tax=D522-D523
Taxation=D524*0.35
Profit after tax=D524-D525
Total Cashflow=D526+D523

Tax on profit=2400000*35%
Net income from sellls of Asstes=D529-D530

Net Present Value=NPV(D532,D535:D544)+D534

I hope my efforts will be fruitful to you...?